# Misaki Kojima    Naoki Nishida

### v1

@@ 3 Reformulation of Inference Rules fo ARSs
We define $\mathsf{DVP}_{+}(\mathcal{A})$ consisting of the rules in Figure˜3.
@@
Note that for an $\mathcal{A}$-runnable set $P$, we have that $\partial_{\mathcal{A}}(P)\neq\emptyset$ for Der, because $P\setminus\mathit{NF\!}_{\mathcal{A}}\neq\emptyset$. Since constrained rewrite rules of an LCTRS may overlap, we allow $P_{i}$ and $P_{j}$ ($i\neq j$) to have overlaps, i.e., $P_{i}\cap P_{j}\neq\emptyset$.
@@
Proof 3.3 (Proof (Sketch)).
The side conditions of rules in $\mathsf{DVP}_{+}(\mathcal{A})$ are non-overlapping.
@@
Proof 3.5 (Proof (Sketch)).
The application of Subsumption in $\mathsf{DVP}(\mathcal{A})$ can be simulated by Subs and Axiom in $\mathsf{DVP}_{+}(\mathcal{A})$ (with $n=1$). The application of Step in $\mathsf{DVP}(\mathcal{A})$ can be simulated by Subs and Der in $\mathsf{DVP}_{+}(\mathcal{A})$ (with $n=1$). Therefore, we have that $\nu\widehat{\mathsf{DVP}(\mathcal{A})}\subseteq\nu\widehat{\mathsf{DVP}_{+}(\mathcal{A})}$.
The application of Axiom in $\mathsf{DVP}_{+}(\mathcal{A})$ can be simulated by Subsumption in $\mathsf{DVP}(\mathcal{A})$. The application of Subs in $\mathsf{DVP}_{+}(\mathcal{A})$ is followed by either Axiom or Der in $\mathsf{DVP}_{+}(\mathcal{A})$. The sequential application of Subs and Axiom can be simulated by Subsumption in $\mathsf{DVP}(\mathcal{A})$. The sequential application of Subs and Der can be simulated by Step in $\mathsf{DVP}(\mathcal{A})$. The application of Der in $\mathsf{DVP}_{+}(\mathcal{A})$—a special case of the successive application of Subs and Der—can be simulated by Step in $\mathsf{DVP}(\mathcal{A})$.
@@
Proof 3.7.
We proceed by contradiction. Assume that $P\cap Q=\emptyset$, $P\cap\mathit{NF\!}_{\mathcal{A}}\neq\emptyset$, and $\mathcal{A}\mathrel{\mathrel{\models^{\forall}_{\mathrm{partial}}}}P\Rightarrow Q$. Then, $P$ includes a normal form $s$ of $\mathcal{A}$: $s\in P\cap\mathit{NF\!}_{\mathcal{A}}$. Since $P\cap Q=\emptyset$, we have that $s\notin Q$. Since $s$ is a normal form of $\mathcal{A}$, $s$ itself is an execution path of $\mathcal{A}$, which starts from $P$. It follows from $\mathcal{A}\mathrel{\mathrel{\models^{\forall}_{\mathrm{partial}}}}P\Rightarrow Q$ that $s\in Q$. This contradicts the fact that $s\notin Q$.
Proposition˜3.6 implies an inference rule for disproof (cf. [KN24jip]).
@@
The set $\mathsf{DVP}_{\bot}(\mathcal{A})$ of inference rules consists of the rules in $\mathsf{DVP}_{+}(\mathcal{A})$ and the following rule for disproof:
@@
By definition, it is clear that for each APR predicate $P\Rightarrow Q$, exactly one rule in $\mathsf{DVP}_{\bot}(\mathcal{A})$ is applicable (cf. [KN24jip]).
@@
Proof 3.10 (Proof (Sketch)).
The side conditions of rules in $\mathsf{DVP}_{\bot}(\mathcal{A})$ are non-overlapping.
@@ 4 Cyclic Proof System for APR Problems of ARSs
In this section, we adapt the cyclic-proof system for partial validity of APR predicates w.r.t. LCTRSs to the abstract APR framework reformulated in Section˜3. To this end, we first revisit rule Circ in Figure˜2, and then define APR pre-proofs for partial validity w.r.t. ARSs.
@@
As in other proof systems, Cut must be powerful but needs a heuristic which may not be easy to automate. As a first step, we do not consider Cut and leave the introduction of Cut to the proof system as future work.
In [KN24jip], a simpler proof system for partial validity w.r.t. LCTRSs has been formulated in the cyclic-proof style [Bro05]. In the proof system, the circularity rule is not explicitly used, but the bud-companion relationship is used instead in proof trees. We reformulate the cyclic-proof system for partial validity w.r.t. ARSs.
@@
  • for all nodes $v\in V$, $\mbox{\sl p}_{j}(v)$ is defined just in case $\mbox{\sl r}(v)$ is a rule with $k$ premises ($1\leq j\leq k)$, and $\frac{\mbox{\sl a}(\mbox{\sl p}_{1}(v))\penalty 10000\ \ldots\penalty 10000\ \mbox{\sl a}(\mbox{\sl p}_{k}(v))}{\mbox{\sl a}(v)}$ is an instance of rule $\mbox{\sl r}(v)$. [^3] If $k=0$, then $\mbox{\sl p}(v)=\epsilon$.
@@
An APR pre-proof for an APR predicate $P\Rightarrow Q$ w.r.t. $\mathcal{A}$ is a pair $(\mathcal{T},\xi)$ of an APR derivation tree $\mathcal{T}=(V,\mbox{\sl a},\mbox{\sl r},\mbox{\sl p})$ (with $v_{0}$ the root node) and a partial mapping $\xi$ from $V_{\mathit{open}}$ to $V\setminus V_{\mathit{open}}$ such that $\mbox{\sl a}(v_{0})=(P\Rightarrow Q)$, and for any open leaf $v$, if $\xi(v)$ is defined, then $\xi(v)$ is a node of $\mathcal{T}$ such that $\mbox{\sl a}(v)=\mbox{\sl a}(\xi(v))$ [^4] From the viewpoint of Cyc, the condition $\mbox{\sl a}(v)=\mbox{\sl a}(\xi(v))$” can be relaxed to $P\subseteq P^{\prime}$”, where $\mbox{\sl a}(v)=(P\Rightarrow Q)$ and $\mbox{\sl a}(v^{\prime})=(P^{\prime}\Rightarrow Q)$. However, as a starting point, we do not use the relaxed condition, and leave it as future work. and $\mbox{\sl r}(\xi(v))=\mbox{\rm{\rm{Der}}}$. An open leaf $v$ with $\xi(v)$ defined is called a bud node of $\mathcal{T}$, and the node $\xi(v)$ is called a companion of $v$. We denote the set of bud nodes in $V$ by $V_{\mathit{bud}}$ ($\subseteq V_{\mathit{open}}$). The APR pre-proof $(\mathcal{T},\xi)$ is said to be closed if any leaf $v\in V$ is either closed or a bud node. The APR pre-proof $(\mathcal{T},\xi)$ is said to be open if $\mathcal{T}$ is not closed, i.e., there exists an open leaf that is not a bud.
@@
Let us consider the ARS $\mathcal{A}_{1}$ in Example˜2.6 again. Let $v_{0},v_{1},\ldots,v_{5}$ nodes such that
@@
Then, $((\{v_{0},v_{1},\ldots,v_{5}\},\mbox{\sl a},\mbox{\sl r},\mbox{\sl p}),\xi)$ is an APR pre-proof for $\{\mathsf{a}\}\Rightarrow\{\mathsf{c},\mathsf{d}\}$ w.r.t. $\mathcal{A}_{1}$. The APR pre-proof is visualized in Figure˜4 (a), where $\dagger$} indicates the bud-companion relationship, i.e., $\xi(v_{3})=v_{0}$.
$\dagger$} $v_{3}:\penalty 10000\ \{\mathsf{a}\}\Rightarrow\{\mathsf{c},\mathsf{d}\}$            (Axiom) $v_{5}:\penalty 10000\ \emptyset\Rightarrow\{\mathsf{c},\mathsf{d}\}$  (Subs) $v_{4}:\penalty 10000\ \{\mathsf{c}\}\Rightarrow\{\mathsf{c},\mathsf{d}\}$  (Der)        $v_{2}:\penalty 10000\ \{\mathsf{b}\}\Rightarrow\{\mathsf{c},\mathsf{d}\}$       (Subs)      $v_{1}:\penalty 10000\ \{\mathsf{b},\mathsf{d}\}\Rightarrow\{\mathsf{c},\mathsf{d}\}$       (Der)      $\dagger$} $v_{0}:\penalty 10000\ \{\mathsf{a}\}\Rightarrow\{\mathsf{c},\mathsf{d}\}{\dagger}$
@@
(a) APR pre-proof (b) the proof graph of (a)
Figure 4: An APR pre-proof and its proof graph for the APR predicate $\{\mathsf{a}\}\Rightarrow\{\mathsf{c},\mathsf{d}\}$.
@@
The proof graph of $((\{v_{0},v_{1},\ldots,v_{5}\},\mbox{\sl a},\mbox{\sl r},\mbox{\sl p}),\xi)$ in Example˜4.3 (i.e., Figure˜8 (a)) is illustrated in Figure˜8 (b).
By definition, nodes of proof graphs have the following properties.
@@
Let $(\mathcal{T},\xi)$ be a closed APR pre-proof, $(V,E)$ the proof graph of $(\mathcal{T},\xi)$, $v,v^{\prime}\in V$, $X\in\{\mbox{\rm{\rm{Axiom}}},\mbox{\rm{\rm{Subs}}},\mbox{\rm{\rm{Der}}},\mbox{\rm{\rm{Dis}}}\}$. Then, all of the following statements hold:
@@
Proof 4.7.
Trivial by definition.
Cyclic proofs satisfy the global trace condition [Bro05]. Roughly speaking, the global trace condition requires every infinite trace—a sequence of atomic formulas in premises sets of sequents—to follow an infinite path of the proof graph, which goes through infinitely many edges corresponding to the application of rules of inductive predicates. In our case, such edges correspond to those from nodes with Der. To make proof graphs implicitly satisfy the global trace condition, we require companions to be nodes with Der; the form of APR predicates can be considered the same as that of sequents, and each APR predicate has exactly one constrained term corresponding to an atomic formula of premises of sequents; viewed in this light, traces correspond to paths of proof graphs of closed APR pre-proofs; every cycle goes through companions infinitely often, and thus every infinite trace follows an infinite path that goes through infinitely many nodes with Der.
@@
Figure 5: An APR pre-proof for $\{\mathsf{a}\}\Rightarrow\{\mathsf{c}\}$, which is an APR disproof
@@
Proof 4.11 (Proof (Sketch)).
The first claim is proved by contradiction, analogously to the proof in [CL18, Theorem 3]. For the second claim, the APR disproof for $P\Rightarrow Q$ has a path from the root to a leaf with $\bot$, which implies a finite execution path that starts from $P$ and does not include any element in $Q$.
@@
    if $\{t\in A\mid\exists s\in P.\ s\to_{\mathcal{A}}^{*}t\}$ is finite, and the emptiness, intersection emptiness, and $\mathcal{A}$-runnability problems for $A$ are decidable, then partial validity of $P\Rightarrow Q$ w.r.t. $\mathcal{A}$ is decidable.
Proof 4.13 (Proof (Sketch)).
For partial validity of $P\Rightarrow Q$, it suffices to consider APR predicates $P^{\prime}\Rightarrow Q$ such that $P^{\prime}\subseteq\{t\in A\mid\exists s\in P.\ s\to_{\mathcal{A}}^{*}t\}$. Since $\{t\in A\mid\exists s\in P.\ s\to_{\mathcal{A}}^{*}t\}$ is finite, there are only finitely many APR predicates to be considered, and thus we can construct a closed APR pre-proof by making generated nodes buds as much as possible. Since the closed APR pre-proof is either an APR proof or disproof, the first claim holds.
For the second claim, since the side conditions of rules in $\mathsf{DVP}_{\bot}(\mathcal{A})$ are decidable by the assumption, we can construct a closed APR pre-proof, which is either an APR proof or disproof. Therefore, the partial validity of $P\Rightarrow Q$ w.r.t. $\mathcal{A}$ is decidable.
@@ 5 Reduction of Safety Properties to APR predicates
Informally speaking, safety properties specify that “something bad never happens” [BK08, Section 3.3.2]. As a safety property w.r.t. a system and a set $E$ of error states, we consider the property that any (possibly non-terminating) execution of the system never reaches any error state in $E$. This kind of safety properties is formulated for ARSs as follows:
@@
This problem is the same as non-reachability from $P$ to $E$. On the other hand, this problem can be reduced to the APR problem [KN23jlamp]. In this section, we revisit the reduction of the non-reachability problem to partial validity of APR predicates w.r.t. LCTRSs, and then reformulate it for ARSs.
@@
    the non-reachability problem w.r.t. the error states $E$ is reduced to the APR predicate $I\Rightarrow\{\mathsf{success}\}$, where $I$ is the set of ground terms representing initial states of a system modeled by $\mathcal{R}$.
@@
We now adapt the above approach to ARSs. Usually, error states are irreducible, and thus we consider only irreducible error states. If an error state to be considered is reducible, then, as in [KN23jlamp], we introduce a fresh constant such as $\mathsf{error}$ above, together with constrained rewrite rules from the error states to $\mathsf{error}$, which is considered a dummy error state in the modified system.
@@
Proof 5.3.
Proof (Sketch) Both the only-if and if parts are proved by contradiction.
@@
Let us consider the program graph in Figure˜6 for Peterson’s mutual exclusion algorithm [BK08, Example 2.25]. Two processes $P_{0},P_{1}$ share Boolean variables $b_{0},b_{1}$ and an integer variable $x$; The Boolean variables $b_{i}$ indicates that $P_{i}$ wants to enter the critical section $\mathsf{crit}_{i}$; The integer variable $x$ stores the identifier of the process that has priority for the critical section at that time. A state of the asynchronous integer transition systems (AITS, for short) with shared variables [BK08] consisting of $\mathit{PG}_{0}$ and $\mathit{PG}_{1}$ are a tuple of a location of $\mathit{PG}_{0}$, a location of $\mathit{PG}_{1}$, and an assignment from variables $b_{0},b_{1},x$ to values. The initial states of the AITS are $\langle\,\mathsf{noncrit}_{0},\mathsf{noncrit}_{1},\mathsf{false},\mathsf{false},m\,\rangle$ with $m\in\{0,1\}$. Let $\mathit{State}_{[}i]$ be the set $\{\mathsf{noncrit}_{i},\mathsf{wait}_{i},\mathsf{crit}_{i}\}$ ($i\in\{0,1\}$), $\mathbb{B}$ the set $\{\mathsf{true},\mathsf{false}\}$, and $\mathcal{S}$ the set of assignments for variables $b_{0},b_{1}:\mbox{\sl{bool}}$ and $x:\mbox{\sl{int}}$: $\mathcal{S}=\{\{b_{0}\mapsto v_{0},b_{1}\mapsto v_{1},x\mapsto m\}\mid v_{0},v_{1}\in\mathbb{B},\penalty 10000\ m\in\mathbb{Z}\}$. We denote a state $\langle\,p_{0},p_{1},\sigma\,\rangle$ by $\langle\,p_{0},p_{1},\sigma(b_{0}),\sigma(b_{1}),\sigma(x)\,\rangle$. The AITS is represented by the following ARS:
@@
Let us consider the race freedom of mutual exclusion—the two processes do not enter their critical sections simultaneously—for the AITS $\mathcal{A}_{2}$. There is no normal form reachable from an initial state. This property will be examined later, together with race freedom. The error states are $\langle\,\mathsf{crit}_{0},\mathsf{crit}_{1},\sigma\,\rangle$ with $\sigma\in\mathcal{S}$. Since all the error states are reducible w.r.t. $\mathcal{A}_{2}$, we introduce a fresh element $\mathsf{error}$ as an error state to be verified, and add the reduction from the error states to $\mathsf{error}$: $\langle\,\mathsf{crit}_{0},\mathsf{crit}_{1},\sigma\,\rangle\to_{\mathcal{A}_{2}}\mathsf{error}$, where $\sigma\in\mathcal{S}$. We let $\mathcal{A}_{2}^{\prime}$ be the modified ARS $(\{\langle\,p_{0},p_{1},\sigma\,\rangle\mid p_{0}\in\mathit{State}_{0},\,p_{1}\in\mathit{State}_{1},\,\sigma\in\mathcal{S}\},\to_{\mathcal{A}_{2}^{\prime}})$, where ${\to_{\mathcal{A}_{2}^{\prime}}}={\to_{\mathcal{A}_{2}}}\cup\{(\langle\,\mathsf{crit}_{0},\mathsf{crit}_{1},\sigma\,\rangle,\mathsf{error})\mid\sigma\in\mathcal{S}\}$. The race freedom is reduced to the following APR predicate:
$\{\langle\,\mathsf{noncrit}_{0},\mathsf{noncrit}_{1},\mathsf{false},\mathsf{false},m\,\rangle\mid m\in\{0,1\}\}\Rightarrow\emptyset$
We obtain an APR proof illustrated in Figure˜7, and thus, by Theorem 4.10, this APR predicate is partially valid w.r.t. $\mathcal{A}_{2}^{\prime}$. Therefore, by Theorem˜5.2, the AITS defined by the program graphs in Figure˜6 is race free and, by Proposition˜2.4 5, there is no normal form reachable from any initial state.
@@
Proof 5.6 (Proof (Sketch)).
Both the only-if and if parts are proved by contradiction.
One may think that non-reachability analysis is more suitable for verification of safety properties. However, APR analysis is essentially very similar to non-reachability analysis. For non-reachability from $P$ to $E$, all states reachable from $P$ are verified, unless an error state reachable from $P$ is detected. On the other hand, during the construction of an APR pre-proof, all states reachable from $P$ are verified, unless a subset $P^{\prime}$ with $P^{\prime}\cap E\neq\emptyset$ is detected.

### v2

@@ 3 Reformulation of Inference Rules for ARSs
We define $\mathsf{DVP}_{+}(\mathcal{A})$ consisting of the inference rules in Figure˜3.
@@
Note that for an $\mathcal{A}$-runnable set $P$, we have that $\partial_{\mathcal{A}}(P)\neq\emptyset$ for Der, because $P\setminus\mathit{NF\!}_{\mathcal{A}}\neq\emptyset$. Note also that $P_{i}$ and $P_{j}$ ($i\neq j$) may overlap, i.e., $P_{i}\cap P_{j}\neq\emptyset$.
@@
Proof 3.5.
We proceed by contradiction. Assume that $P\cap Q=\emptyset$, $P\cap\mathit{NF\!}_{\mathcal{A}}\neq\emptyset$, and $\mathcal{A}\mathrel{\mathrel{\models^{\forall}_{\mathrm{partial}}}}P\Rightarrow Q$. Then, there exists a normal form $s\in P\cap\mathit{NF\!}_{\mathcal{A}}$ such that $s\notin Q$. It follows from $\mathcal{A}\mathrel{\mathrel{\models^{\forall}_{\mathrm{partial}}}}P\Rightarrow Q$ that $s\in Q$. This contradicts the fact that $s\notin Q$.
Proposition˜3.4 implies an inference rule for disproof (cf. [KN24jip]).
@@
We define $\mathsf{DVP}_{\bot}(\mathcal{A})$ consisting of the inference rules in $\mathsf{DVP}_{+}(\mathcal{A})$ and the following rule for disproof:
@@
By definition, for each APR predicate, exactly one rule in $\mathsf{DVP}_{\bot}(\mathcal{A})$ is applicable (cf. [KN24jip]).
@@ 4 Cyclic Proof System for APR Problems of ARSs
In this section, we adapt the cyclic-proof system for partial validity of APR predicates w.r.t. LCTRSs to the abstract APR framework reformulated in Section˜3. To this end, we first revisit Circ in Figure˜2, and then define APR pre-proofs for partial validity w.r.t. ARSs.
@@
As in other proof systems, Cut must be powerful but needs a heuristic for automation. As a first step, we leave the introduction of Cut to the proof system below as future work.
In [KN24jip], a simpler proof system for partial validity w.r.t. LCTRSs has been formulated in the cyclic-proof style [Bro05]. In the proof system, the circularity rule is not explicitly used, but the bud-companion relationship is used instead in proof trees. We formulate the cyclic-proof system for partial validity w.r.t. ARSs.
@@
  • for all nodes $v\in V$, $\mbox{\sl p}_{j}(v)$ is defined just in case $\mbox{\sl r}(v)$ is a rule with $k$ premises ($1\leq j\leq k)$, and $\frac{\mbox{\sl a}(\mbox{\sl p}_{1}(v))\penalty 10000\ \ldots\penalty 10000\ \mbox{\sl a}(\mbox{\sl p}_{k}(v))}{\mbox{\sl a}(v)}$ is an instance of rule $\mbox{\sl r}(v)$, where if $k=0$, then $\mbox{\sl p}(v)=\epsilon$.
@@
An APR pre-proof for an APR predicate $P\Rightarrow Q$ w.r.t. $\mathcal{A}$ is a pair $(\mathcal{T},\xi)$ of an APR derivation tree $\mathcal{T}=(V,\mbox{\sl a},\mbox{\sl r},\mbox{\sl p})$ (with $v_{0}$ the root node) and a partial mapping $\xi$ from $V_{\mathit{open}}$ to $V\setminus V_{\mathit{open}}$ such that $\mbox{\sl a}(v_{0})=(P\Rightarrow Q)$, and for any open leaf $v$, if $\xi(v)$ is defined, then $\xi(v)$ is a node of $\mathcal{T}$ such that $\mbox{\sl a}(v)=\mbox{\sl a}(\xi(v))$ [^3] From the viewpoint of Cyc, the condition $\mbox{\sl a}(v)=\mbox{\sl a}(\xi(v))$” can be relaxed to $\mbox{\sl a}_{L}(v)\subseteq\mbox{\sl a}_{L}(\xi(v))$”. However, as a starting point, we do not use the relaxed condition, and leave it as future work. and $\mbox{\sl r}(\xi(v))=\mbox{\rm{\rm{Der}}}$. An open leaf $v$ with $\xi(v)$ defined is called a bud node of $\mathcal{T}$, and the node $\xi(v)$ is called a companion of $v$. We denote the set of bud nodes in $V$ by $V_{\mathit{bud}}$ ($\subseteq V_{\mathit{open}}$). The APR pre-proof $(\mathcal{T},\xi)$ is said to be closed if any leaf $v\in V$ is either closed or a bud node. The APR pre-proof $(\mathcal{T},\xi)$ is said to be open if $\mathcal{T}$ is not closed, i.e., there exists an open leaf that is not a bud.
@@
Let us consider the ARS $\mathcal{A}_{1}$ in Example˜2.5 again. Let $v_{0},v_{1},\ldots,v_{5}$ nodes such that
@@
Then, $((\{v_{0},v_{1},\ldots,v_{5}\},\mbox{\sl a},\mbox{\sl r},\mbox{\sl p}),\xi)$ is an APR pre-proof for $\{\mathsf{a}\}\Rightarrow\{\mathsf{c},\mathsf{d}\}$ w.r.t. $\mathcal{A}_{1}$, which is visualized in Figure˜4 (a).
$\dagger$} $v_{3}:\penalty 10000\ \{\mathsf{a}\}\Rightarrow\{\mathsf{c},\mathsf{d}\}$            (Axiom) $v_{5}:\penalty 10000\ \emptyset\Rightarrow\{\mathsf{c},\mathsf{d}\}$  (Subs) $v_{4}:\penalty 10000\ \{\mathsf{c}\}\Rightarrow\{\mathsf{c},\mathsf{d}\}$  (Der)        $v_{2}:\penalty 10000\ \{\mathsf{b}\}\Rightarrow\{\mathsf{c},\mathsf{d}\}$       (Subs)      $v_{1}:\penalty 10000\ \{\mathsf{b},\mathsf{d}\}\Rightarrow\{\mathsf{c},\mathsf{d}\}$       (Der)       $\dagger$} $v_{0}:\penalty 10000\ \{\mathsf{a}\}\Rightarrow\{\mathsf{c},\mathsf{d}\}$
@@
(a) APR pre-proof (b) proof graph
Figure 4: An APR pre-proof and its proof graph for the APR predicate $\{\mathsf{a}\}\Rightarrow\{\mathsf{c},\mathsf{d}\}$, where $\dagger$} indicates the bud-companion relationship.
@@
The proof graph of $((\{v_{0},v_{1},\ldots,v_{5}\},\mbox{\sl a},\mbox{\sl r},\mbox{\sl p}),\xi)$ in Example˜4.3 (i.e., Figure˜4 (a)) is illustrated in Figure˜4 (b).
By definition, nodes of proof graphs trivially have the following properties.
@@
Let $(\mathcal{T},\xi)$ be a closed APR pre-proof, $(V,E)$ the proof graph of $(\mathcal{T},\xi)$, $v,v^{\prime}\in V$, and $X\in\{\mbox{\rm{\rm{Axiom}}},\mbox{\rm{\rm{Subs}}},\mbox{\rm{\rm{Der}}},\mbox{\rm{\rm{Dis}}}\}$. Then, all of the following statements hold:
@@
Cyclic proofs satisfy the global trace condition [Bro05]. Roughly speaking, the global trace condition requires every infinite trace—a sequence of atomic formulas in premises sets of sequents—to follow an infinite path of the proof graph, which goes through infinitely many edges corresponding to the application of rules of inductive predicates. In our case, such edges correspond to those from nodes with Der. To make proof graphs implicitly satisfy the global trace condition, we require companions to be nodes with Der; the form of APR predicates can be considered the same as that of sequents, and each APR predicate has exactly one set corresponding to an atomic formula of premises of sequents; viewed in this light, traces correspond to paths of proof graphs of closed APR pre-proofs; every cycle goes through companions infinitely often, and thus every infinite trace follows an infinite path that goes through infinitely many nodes with Der.
@@
Figure 5: An APR pre-proof for $\{\mathsf{a}\}\Rightarrow\{\mathsf{c}\}$, which is an APR disproof.
@@
    if $\{t\in A\mid\exists s\in P.\ s\to_{\mathcal{A}}^{*}t\}$ is finite and the emptiness, intersection emptiness, and $\mathcal{A}$-runnability problems for $A$ are decidable, then partial validity of $P\Rightarrow Q$ w.r.t. $\mathcal{A}$ is decidable.
@@ 5 Reduction of Safety Properties to APR predicates
Informally speaking, safety properties specify that “something bad never happens” [BK08, Section 3.3.2]. As a safety property w.r.t. a system and a set $E$ of error states, we consider the property that any (possibly non-terminating) execution of the system never reaches any error state in $E$. This kind of safety properties is formulated as a problem for ARSs as follows:
@@
This problem is the same as non-reachability from $P$ to $E$ w.r.t. $\mathcal{A}$. On the other hand, this problem can be reduced to the APR problem [KN23jlamp]. In this section, we revisit the reduction of the non-reachability problem to partial validity of APR predicates w.r.t. LCTRSs, and then reformulate it for ARSs.
@@
    the non-reachability problem w.r.t. the error states $E$ is reduced to the APR predicate $I\Rightarrow\{\mathsf{success}\}$, where $I$ is the set of ground terms for initial states of a system.
@@
We now adapt the above approach to ARSs. Usually, error states are irreducible, and thus we consider only irreducible error states. If an error state is reducible, then, as in the approach above for LCTRSs, we introduce a fresh constant such as $\mathsf{error}$, together with the additional reduction from the error states to $\mathsf{error}$, which is considered a dummy error state in the modified system.
@@
Let us consider the program graphs $\mathit{PG}_{0},\mathit{PG}_{1}$ in Figure˜6 for Peterson’s mutual exclusion algorithm [BK08, Example 2.25]. Two processes $P_{0},P_{1}$ specified by $\mathit{PG}_{0},\mathit{PG}_{1}$, respectively, share Boolean variables $b_{0},b_{1}$ and an integer variable $x$; The Boolean variables $b_{i}$ indicates that $P_{i}$ wants to enter the critical section $\mathsf{crit}_{i}$; The integer variable $x$ stores the identifier of the process that has priority for the critical section at that time. A state of the asynchronous integer transition systems (AITS, for short) with shared variables [BK08] consisting of $\mathit{PG}_{0}$ and $\mathit{PG}_{1}$ are a tuple of a location of $\mathit{PG}_{0}$, a location of $\mathit{PG}_{1}$, and an assignment from variables $b_{0},b_{1},x$ to values. The initial states of the AITS are $\langle\,\mathsf{noncrit}_{0},\mathsf{noncrit}_{1},\mathsf{false},\mathsf{false},m\,\rangle$ with $m\in\{0,1\}$. Let $\mathit{State}_{i}$ be the set $\{\mathsf{noncrit}_{i},\mathsf{wait}_{i},\mathsf{crit}_{i}\}$ ($i\in\{0,1\}$), $\mathbb{B}$ the set $\{\mathsf{true},\mathsf{false}\}$, and $\mathcal{S}$ the set of assignments for variables $b_{0},b_{1}:\mbox{\sl{bool}}$ and $x:\mbox{\sl{int}}$: $\mathcal{S}=\{\{b_{0}\mapsto v_{0},b_{1}\mapsto v_{1},x\mapsto m\}\mid v_{0},v_{1}\in\mathbb{B},\penalty 10000\ m\in\mathbb{Z}\}$. We denote a state $\langle\,p_{0},p_{1},\sigma\,\rangle$ by $\langle\,p_{0},p_{1},\sigma(b_{0}),\sigma(b_{1}),\sigma(x)\,\rangle$. The AITS consisting of $P_{0}$ and $P_{1}$ is represented by the following ARS:
@@
Let us consider the race freedom of mutual exclusion—the two processes do not enter their critical sections simultaneously—for the AITS $\mathcal{A}_{2}$. There is no normal form reachable from an initial state. This property will be examined later, together with race freedom. The error states are $\langle\,\mathsf{crit}_{0},\mathsf{crit}_{1},\sigma\,\rangle$ with $\sigma\in\mathcal{S}$. Since all the error states are reducible w.r.t. $\mathcal{A}_{2}$, we introduce a fresh element $\mathsf{error}$ as a dummy error state to be verified, and add the reduction from the original error states to $\mathsf{error}$: $\langle\,\mathsf{crit}_{0},\mathsf{crit}_{1},\sigma\,\rangle\to_{\mathcal{A}_{2}}\mathsf{error}$, where $\sigma\in\mathcal{S}$. We let $\mathcal{A}_{2}^{\prime}$ be the modified ARS $(\{\langle\,p_{0},p_{1},\sigma\,\rangle\mid p_{0}\in\mathit{State}_{0},\,p_{1}\in\mathit{State}_{1},\,\sigma\in\mathcal{S}\},\to_{\mathcal{A}_{2}^{\prime}})$, where ${\to_{\mathcal{A}_{2}^{\prime}}}={\to_{\mathcal{A}_{2}}}\cup\{(\langle\,\mathsf{crit}_{0},\mathsf{crit}_{1},\sigma\,\rangle,\mathsf{error})\mid\sigma\in\mathcal{S}\}$. The race freedom is reduced to the APR predicate $(1)\penalty 10000\ \{\langle\,\mathsf{noncrit}_{0},\mathsf{noncrit}_{1},\mathsf{false},\mathsf{false},m\,\rangle\mid m\in\{0,1\}\}\Rightarrow\emptyset$. We obtain an APR proof illustrated in Figure˜7, and thus, by Theorem 4.9, the APR predicate (1) is partially valid w.r.t. $\mathcal{A}_{2}^{\prime}$. Therefore, by Theorem˜5.2, the AITS defined by the program graphs in Figure˜6 is race free and, by Proposition˜2.4 5, there is no normal form reachable from any initial state.
@@
One may think that non-reachability analysis is more suitable for safety properties. However, APR analysis is essentially very similar to non-reachability analysis; for non-reachability from $P$ to $E$, all states reachable from $P$ are examined, unless an error state reachable from $P$ is detected; during the construction of an APR pre-proof, all states reachable from $P$ are examined, unless a subset $P^{\prime}$ with $P^{\prime}\cap E\neq\emptyset$ is detected.